OpenStudy (anonymous):

How can you make a dollar using 10 coins (must use 10) and only using nickels quarters pennies and half dollars not dimes!

jimthompson5910 (jim_thompson5910):

Here are the 6 ways I found 10 dimes ... 10 coins total 10*0.10 = 1.00 3 nickels + 6 dimes + 1 quarter ... 10 coins total 3*0.05 + 6*0.10+1*0.25 = 1.00 6 nickels + 2 dimes + 2 quarters ... 10 coins total 6*0.05 + 2*0.10+2*0.25 = 1.00 8 nickels + 1 dime + 1 half dollar ... 10 coins total 8*0.05 + 1*0.10+1*0.5 = 1.00 5 pennies + 2 dimes + 3 quarters ... 10 coins total 5*0.01 + 2*0.10 + 3*0.25 = 1.00 5 pennies + 2 nickels + 1 dime + 1 quarter + 1 half dollar ... 10 coins total 5*0.01 + 2*0.05 + 1*0.10 + 1*0.25 + 1*0.50 = 1.00 All of them require at least one dime. So it looks to be impossible to do it without dimes. I might be missing one possibility though.

OpenStudy (anonymous):

thx :D

jimthompson5910 (jim_thompson5910):

np